# Algorithmic and High-Frequency Trading Problems for Semi-Markov and Hawkes Jump-Diffusion Models

### v1

@@ Abstract
Algorithmic and High-Frequency trading (HFT) has become one of the main ways to complete transactions in many of today’s major financial markets, with these transactions taking place inside what is called the limit order book (LOB). Developing sophisticated trading algorithms that accurately mimic LOB data is therefore a major topic in this area. In recent times, it has been proven that LOB data often follows non-Markovian dynamics, thus, we believe these models more accurately describe how the LOB would evolve. In this paper, we consider acquisition and liquidation problems for semi-Markov and Hawkes jump-diffusion models. We begin by developing jump-diffusion models to capture these dynamics and then proceed to use diffusion approximations for the jump parts. The optimal solutions to these trading problems are formulated under the stochastic optimal control framework and via numerical methods. Strategy simulations for the acquisition and liquidation problems are considered as well, where we show sample price paths for our price processes, average traded prices, inventory and trading speed paths. This analysis gives a general picture of how one could analyse how these strategies could perform under our more general price processes.
@@ 1 Introduction
The focus of this paper will involve formulating asset price processes of the jump-diffusion type, through a diffusion approximation for the jump parts. In general, we aim to develop stochastic midprice processes, which are referred to as the “fundamental price” in Cartea et al., (2015), that portray an accurate representation of how the value of a financial asset evolves. In particular, we focus on assets where the majority of transactions take place within the LOB, thus the evolution of the price here would generally represent the value of these assets. As new information about these financial assets arrives into the market or certain transactions take place (the larger the more effect they tend to have), the fundamental price then evolves. This is often modelled as a combination of noise via the increments of Brownian motion along with some added drift that may be occurring, often based on certain traders actions. If the new information entering the market is very sudden and ground breaking, or the transactions taking place are abnormally large, big jumps can instantly occur in the fundamental price. In this case, a jump model would be necessary to accurately model the midprice dynamics.
We will begin by defining a general version of the stochastic midprice processes used in Cartea et al., (2015), as they regularly use this midprice process to solve SOC trading problems. This midprice process can be defined as follows,
@@
The models in this paper aim to extend and improve the stochastic midprice processes used for modelling LOB dynamics, under the SOC framework. High frequency LOB data often exhibits points of discontinuity i.e., jumps to different price levels, which a pure diffusion model would ignore. These jumps in the price process can occur for a number reasons, such as a random news event entering the market which erodes liquidity at nearby price levels, or a random large buyer/seller enters the market with a high urgency to complete a trade. Thus, it’s paramount for a price process to be able to model these dynamics in order to better specify financial asset pricing dynamics. Recent studies have shown that jump-diffusion pricing dynamics best represent the majority of the LOB data in financial markets. Upon examining the data, one would notice that it regularly follows diffusion dynamics, but the data often experiences jumps as well. As well as exhibiting jump dynamics, LOB data, in general, tends to follow non-markovian dynamics as shown by many extensive studies on the data. The empirical results in He and Swishchuk, (2019), Swishchuk et al., (2019) and Swishchuk and Huffman, (2020) shows that higher numbers of state dependent orders are often more accurate in modelling LOB midprice dynamics versus the mathematical assumption of an infinitesimal tick size in a general GBM model. Similar evidence can be found in many more works and some examples, in no particular order, include Cartea et al., 2018a , Cartea et al., 2018b , Drame, (2019), Mäkinen et al., (2019), Sfendourakis and Toke, (2021), Sjogren and DeLise, (2021), and many more. Upon examining the LOB data, these papers all build models with the underlying assumption that jumps in price occur.
@@
Thus, our aim in this paper is to add more versatility to the pricing models used to formulate and solve algorithmic trading problems, like in Cartea et al., (2015), where they mostly use midprice processes based on a pure arithmetic diffusion model as shown in equation (1). Therefore, our models are new and more general for both the cases considered in: i) Cartea et al., (2015) for diffusion case only with drift, see (1), and ii) in Roldan Contreras and Swishchuk, (2022) and Roldan Contreras, (2023), where the pricing models only considered jumps by applying diffusion approximations for the Semi-Markov LOB models in Swishchuk and Vadori, (2017) and the General Compound Hawkes Process (GCHP) LOB models in Swishchuk and Huffman, (2020). We aim to extend both cases, the first one with only diffusion and the second one with only jumps, by formulating more general price processes to account for these dynamics.
@@
The paper is organized as follows. Section 2 introduces jump-diffusion models in HFT, namely, semi-Markov and Hawkes process ones. Section 3 studies two stochastic optimal control trading problems in algorithmic and HFT, namely an acquisition and liquidation problem, under our new jump-diffusion models. Numerical solutions to the acquisition and liquidation problems are considered in Section 4. Strategy simulations for the acquisition and liquidation problems are presented in Section 5. Lastly, a conclusion and future recommendations are discussed in Section 6.
@@ 2.1 Jumps as a function of a Semi-Markov Process
To first give some background, previous models, such as in Cont and De Larrard, (2013), suggest that for modelling the LOB, one can use a straightforward stochastic model for the dynamics of a LOB, in which the arrivals of MOs, LOs, and order cancellations are described in terms of a Markovian queueing system. In Swishchuk and Vadori, (2017), this work was extended to reflect the fact that an arbitrary distribution for the inter-arrival times of orders inside the LOB tend to be non-exponential. They portray how the nature of a new book event and its corresponding inter-arrival times depend on the nature of the previous book event. This contradicts the Markovian model in Cont and De Larrard, (2013) and the use of a Poisson and Compound Poisson process to model the jump part of the stock price process. To improve the model to reflect this non-Markovian property, Fodra and Pham, 2015b uses a Markov Renewal process, also known as a Semi-Markov process, with this paper focusing on this from a market microstructure perspective. In Fodra and Pham, 2015a , they study an optimal high frequency trading problem where the price process is driven by a Markov Renewal Process.
Semi-Markov processes are formed from a wide array of stochastic processes. The main advantage over a regular Markov process is that it allows the use of an arbitrary waiting time distribution for modelling the time until a transition occurs out of one state to another i.e., waiting times do not have to be modelled using an exponential distribution. Past work first developed a General Semi-Markov model (in Swishchuk et al., (2017)) for modelling the dynamics of LOBs and then Swishchuk and Vadori, (2017) developed an approximation to the Semi-Markov process as a diffusion model. In Roldan Contreras, (2023), this led to an additional drift term in the stock price process when they solved SOC algorithmic trading problems. In this paper, we first formulate a model for the case where the price process follows the regular diffusion model from equation (1) plus a jump part, i.e., a jump-diffusion model, as follows:
@@
In equation (2), $W_{t}$, $\sigma$ and $g$ are as defined before, $X_{k}$ is an ergodic Markov chain representing two types of consecutive price increments taking values ±$\pm\delta$ and $N(t)$ is the counting process for these price changes. This price process now includes an additional jump term, $\sum_{k=1}^{N(t)}X_{k}$, which is not included in equation (1). Thus, we can gain more insight into how a SOC algorithmic trading problem would perform using a price process that includes the dynamics of this jump part.
@@
where the SM index refers to the fact that this coefficient originates from a Semi-Markov process. Later on, for the Hawkes process case, we will introduce a similar coefficient, $\eta_{HP}$, thus we index this coefficient to avoid confusion later on. The normalized prices $s^{*}$ is defined as $s^{*}:=\delta(2\pi^{*}-1)$, where $\pi^{*}$ is a long-run probability. For a full proof to show that the diffusion part of this renormalized price process satisfies a weak convergence in the Skorokhod topology, see Swishchuk and Vadori, (2017).
Following along a similar fashion, $\varsigma_{SM}=\frac{\sigma^{*}}{\sqrt{\tau}}$ is another new coefficient representing the Semi-Markov diffusion approximation for the jump part of equation (2). Here, $\sigma^{*}$ is a constant depending on the ergodic and transition probabilities of the Markov chain $X_{k}$ and $\tau$ refers to the interarrival time of the jumps. Then for our two state Markov chain, $\sigma^{*}$ can be defined, as in Swishchuk and Vadori, (2017), as,
$\sigma^{*}=\sqrt{4\delta^{2}\left(\frac{1-p_{cont}^{\text{'}}+\pi^{*}(p_{cont}^{\text{'}}-p_{cont})}{(p_{cont}+p_{cont}^{\text{'}}-2)^{2}}\right)},$ (5)
@@ 2.2 Jumps as a function of a Hawkes Process
where $g$ and the diffusion term are as defined before in section 2.1. In equation (7), $a$ maps the current state to a price movement, $X_{k}$ is again a ergodic Markov chain with ergodic probabilities $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$, where the current state is computed using information from the previous state. Then using the diffusion approximation for the jump part in Swishchuk and Huffman, (2020), this becomes,
$\begin{split}S_{t}&=S_{0}\pm(g(\nu_{t})\eta_{HP})t+\sqrt{\sigma^{2}+\bar{\sigma}^{2}_{HP}+\left(\sigma^{*}\sqrt{\frac{\lambda}{1-\mu}}\right)^{2}}W_{t},\ S_{0}=S\\ &=S_{0}\pm(g(\nu_{t})\eta_{HP})t+\sqrt{\sigma^{2}+\bar{\sigma}^{2}_{HP}+\varsigma_{HP}^{2}}W_{t},\ S_{0}=S.\end{split}$ (9)
Here, the drift coefficient can be defined, as in Roldan Contreras, (2023), as,
@@
where $a^{*}$ is a constant depending on the transition probabilities of the Markov chain. This can be defined as in Swishchuk and Huffman, (2020) as $a^{*}:=\sum_{i\in X}\pi_{i}^{*}a(i)$, where $\pi^{*}$ are the ergodic probabilities of the Markov chain $X$. $\lambda$ is again the background intensity and $\mu$ refers to the output of the response/excitation function. Here, $\varsigma_{HP}=\sigma^{*}\sqrt{\frac{\lambda}{1-\hat{\mu}}}$. $\sigma^{*}$ is again a constant depending on the ergodic probabilities of the Markov chain $X_{k}$. Since this kind of Markov chain has two states, where $X_{k}\in\{-\delta,\delta\}$, $\sigma^{*}$ can be defined, as shown by corollary 3 in Swishchuk and Huffman, (2020), as

### v2

@@ Abstract
This paper introduces a jump-diffusion pricing model specifically designed for algorithmic trading and high-frequency trading (HFT). The model incorporates independent jump and diffusion processes, providing a more precise representation of the limit order book (LOB) dynamics within a scaling-limit framework. Given that algorithmic and HFT strategies now dominate major financial markets, accurately modeling LOB dynamics is crucial for developing effective trading algorithms. Recent research has shown that LOB data often exhibit non-Markovian properties, reinforcing the need for models that better capture its evolution. In this paper, we address acquisition and liquidation problems under more general compound semi-Markov and Hawkes jump-diffusion models. We first develop jump-diffusion frameworks to capture these dynamics and then apply diffusion approximations to the jump components so that robust solutions can be given. Optimal trading strategies are formulated using stochastic optimal control (SOC) and solved numerically. Finally, we present strategy simulations analyzing price paths, inventory evolution, trading speed, and average execution prices. This study provides insights into how these models can improve execution strategies under more general price dynamics.
@@ 1 Introduction
The focus of this paper will involve formulating asset price processes of the jump-diffusion type, through a diffusion approximation for the jump parts. In general, we aim to develop stochastic midprice processes, which are referred to as the “fundamental price” in Cartea et al., (2015), that portray an accurate representation of how the value of a financial asset evolves. In particular, we focus on assets where the majority of transactions take place within the LOB, thus the evolution of the price here would generally represent the value of these assets. As new information about these financial assets arrives into the market or certain transactions take place (the larger the more effect they tend to have), the fundamental price then evolves. This is often modeled as a combination of noise via the increments of Brownian motion along with some added drift that may be occurring, often based on certain traders actions. If the new information entering the market is very sudden and ground breaking, or the transactions taking place are abnormally large, big jumps can instantly occur in the fundamental price. In this case, a jump model would be necessary to accurately model the midprice dynamics.
We will begin by defining a general version of the stochastic midprice processes used in Cartea et al., (2015), as they regularly use this midprice process to solve Stochastic Optimal Control (SOC) trading problems. This midprice process can be defined as follows,
@@
The models in this paper aim to extend and improve the stochastic midprice processes used for modeling LOB dynamics, under the SOC framework. High frequency LOB data often exhibit points of discontinuity i.e., jumps to different price levels, which a pure diffusion model would ignore. Jumps in the price process can occur for various reasons, such as unexpected news events that erode liquidity at nearby price levels or the sudden arrival of a large buyer or seller with high urgency to execute a trade. Thus, it is essential for a price process to incorporate these dynamics to more precisely characterize financial asset pricing behavior.
@@
Recent studies such as Cont and De Larrard, (2012), Swishchuk and Vadori, (2017), and Kreher and Milbradt, (2023) have shown that jump-diffusion pricing dynamics is a good approximation technique for LOB data in financial markets. Upon visualizing the data, as can be done by looking at Figure 1, one would notice that it appears to closely resemble diffusion dynamics, but at some points the data also experience sharp jumps. In addition to exhibiting jump dynamics, LOB data, in general, tend to follow non-Markovian dynamics as shown by many extensive studies on the data. The empirical results in He and Swishchuk, (2019), Swishchuk et al., (2019) and Swishchuk and Huffman, (2020) show that higher numbers of state-dependent orders are often more accurate in modeling LOB midprice dynamics versus the mathematical assumption of an infinitesimal tick size in a general GBM model. Similar evidence can be found in many more works and some examples, in no particular order, include Cartea et al., 2018a , Cartea et al., 2018b , Drame, (2019), Mäkinen et al., (2019), Sfendourakis and Toke, (2021), Sjogren and DeLise, (2021), and many more. Upon examining the LOB data, these papers all build models with the underlying assumption that jumps in price occur.
Figure 1: Midprice evolution in Microsoft’s stock on June 21, 2012, from 9:30 AM to 10:30 AM Eastern Time (ET).
@@
Thus, our aim in this paper is to add more versatility to the pricing models used to formulate and solve algorithmic trading problems, as seen in Cartea et al., (2015), where they mostly use midprice processes based on a pure arithmetic diffusion model, as shown in Equation (1). Therefore, our models are new and more general for both the cases considered in: i) Cartea et al., (2015) for diffusion case only with drift, see (1), and ii) in Roldan Contreras and Swishchuk, (2022) and Roldan Contreras, (2023), where the pricing models only considered jumps by applying diffusion approximations for the Semi-Markov LOB models in Swishchuk and Vadori, (2017) and the General Compound Hawkes Process (GCHP) LOB models in Swishchuk and Huffman, (2020). We aim to extend both cases, the first one with only diffusion and the second one with only jumps, by formulating more general price processes to account for these dynamics.
The main contributions of this paper can be summarized as follows.
  1. 1.
    We propose a jump-diffusion pricing model tailored for algorithmic and HFT, integrating independent jump and diffusion processes that more accurately capture LOB dynamics under a scaling-limit approximation. Our framework uniquely allows the jump dynamics to be modeled using either a Semi-Markov or Hawkes process, combined with an independent Brownian motion component. While a pure jump model based on Semi-Markov and Hawkes dynamics in algorithmic and HFT settings is not new, the incorporation of standardized non-Markovian jump components alongside a conventional arithmetic Brownian motion via Jump-Diffusion models in practical applications represents a significant advancement.
  2. 2.
    To facilitate more robust solutions for standard algorithmic and high-frequency trading (HFT) models in limit order book modeling, where transactions occur every millisecond, we provide approximations for the jump components in our jump-diffusion pricing models. In applications, one typically considers a longer time scale, $tn$ (minutes, hours, etc.), rather than $t$ (milliseconds), allowing sufficient time for solving standard algorithmic and HFT trading problems. In this context, limit theorems such as the law of large numbers (averaging) and the functional central limit theorem (diffusion approximation), which apply in the limit as $n\to\infty$ while $t$ remains fixed, become particularly useful.
  3. 3.
    Using real LOB data, we used calibrated model parameters and conducted an in-depth numerical analysis of both acquisition and liquidation problems under our proposed pricing dynamics. This was followed by extensive simulations across various scenarios where the jump component varies significantly, capturing different market regimes. Our analysis demonstrates how the solutions and simulation outcomes can vary considerably in response to these shifting jump dynamics.
@@
The paper is organized as follows. Section 2 introduces jump-diffusion models in HFT, namely, semi-Markov and Hawkes process ones. Section 3 studies two stochastic optimal control trading problems in algorithmic and HFT, namely an acquisition and liquidation problem, under our new jump-diffusion models. Numerical solutions to these acquisition and liquidation problems are considered in Section 4. Strategy simulations for these acquisition and liquidation problems are presented in Section 5. Lastly, our conclusions and future recommendations are discussed in Section 6.
@@ 2 Semi-Markov and Hawkes Jump-Diffusion Models in HFT
This Section is devoted to the modeling of Jump-Diffusion pricing models, whereby the jump parts are modeled via Semi-Markov and Hawkes process dynamics. In Cartea et al., (2015), Markovian jump-diffusion pricing models are mentioned with a diffusion part $W(t)$ and a counting process $N(t)$. However, these models lack robustness when applied to algorithmic and HFT problems, making them unsuitable for practical implementation. Moreover, no existing approach has been proposed for addressing more general jump-diffusion frameworks, such as those involving Compound Semi-Markov and Compound Hawkes processes, in real-world applications. Subsequently, here we propose to approximate the jump parts of these price processes to obtain a diffusion approximation, allowing our more general compound Semi-Markov or compound Hawkes processes to be included in trading problem solutions. Thus, this work recognizes that trading agents often face multiple risk factors, as also identified in Cartea et al., (2015), and proposes a more comprehensive modeling framework to effectively account for these factors in practical applications.
An important consideration with this type of pricing model, as seen in Equation (1) from Cartea et al., (2015) and later in our own models, is its potential drawback: the possibility of negative prices, a limitation that can be further exacerbated by the presence of jumps. To address this, the pricing model should ensure that initial prices are much higher than the potential jump tick sizes, where an extreme amount of jump movements in one direction would have to occur for prices to go negative. As we will show in our simulations later, the price process remains well above negative values, even under extreme scenarios. To fully mitigate this drawback and eliminate any ambiguity, this issue could be addressed by adopting an exponential jump-diffusion pricing model. A similar approach was formulated for the pure jump case in Guo et al., (2022) within a limit order book (LOB) framework.
The remainder of this Section will proceed as follows. In Section 2.1, we introduce our jump-diffusion model under our Semi-Markov approximation approach (via Law of Large Numbers) for the jump part. In Section 2.2, we similarly introduce our jump-diffusion model but under the Hawkes process approximation approach (via a functional central limit theorem) for the jump part.
@@ 2.1 Jumps as a function of a Semi-Markov Process
To first give some background, previous models, such as in Cont and De Larrard, (2013), suggest that for modeling the LOB, one can use a straightforward stochastic model for the dynamics of an LOB, in which the arrivals of market orders, limit orders, and cancellation orders are described in terms of a Markovian queuing system. In Swishchuk and Vadori, (2017), this work was extended to reflect the fact that an arbitrary distribution for the inter-arrival times of orders inside the LOB tend to be non-exponential. They portray how the nature of a new book event and its corresponding inter-arrival times depend on the nature of the previous book event. This contradicts the Markovian model in Cont and De Larrard, (2013) and the use of a Poisson and Compound Poisson process to model the jump part of the stock price process. To improve the model to reflect this non-Markovian property, Fodra and Pham, 2015b uses a Markov Renewal process, also known as a Semi-Markov process, with this paper focusing on this from a market microstructure perspective. In Fodra and Pham, 2015a , they study an optimal high frequency trading problem where the price process is driven by a Markov Renewal Process.
Semi-Markov processes are formed from a wide array of stochastic processes. The main advantage over a regular Markov process is that it allows the use of an arbitrary waiting time distribution for modeling the time until a transition occurs out of one state to another i.e., waiting times do not have to be modeled using an exponential distribution. Past work first developed a General Semi-Markov model (in Swishchuk et al., (2017)) for modeling the dynamics of LOBs and then Swishchuk and Vadori, (2017) developed an approximation to the Semi-Markov process as a diffusion model. In Roldan Contreras, (2023), this led to an additional drift term in the stock price process when they solved SOC algorithmic trading problems. In this paper, we first formulate a model for the case where the price process follows the regular diffusion model from Equation (1) plus a jump part, i.e., a jump-diffusion model, as follows:
@@
In Equation (2), $W_{t}$, $\sigma$ and $g$ are as defined before, $X_{k}$ is an ergodic Markov chain representing two types of consecutive price increments taking values ±$\pm\delta$ and $N(t)$ represents the number of renewals (or arrivals) that have occurred up to time $t$ i.e., $N(t)$ is the counting process for the price changes. This price process now includes an additional jump term, $\sum_{k=1}^{N(t)}X_{k}$, which is not included in Equation (1). $N(t)$ is a counting process independent of $X_{k}$ and $W_{t}$ i.e., all sources of randomness are assumed to be independent. This means that the number of price changes, $N(t)$, does not depend on the tick size described by $X_{k}$. Thus, we can gain more insight into how an SOC algorithmic trading problem would perform using a price process that includes the dynamics of this jump part.
@@
where the SM index refers to the fact that this coefficient originates from a Semi-Markov process. For the Hawkes process case, we will introduce a similar coefficient, $\eta_{HP}$, thus, we index this coefficient to avoid confusion later on. Here, $m_{\tau}=\sum_{i\in\{-\delta,\delta\}}\pi^{*}(i)m(i)$ and the normalized prices $s^{*}$ is defined as $s^{*}:=\delta(2\pi^{*}-1)$, where $\pi^{*}$ is a long-run probability. In the diffusion approximation, we use the scaling limit, $\lim_{n\to\infty}$, for $S_{tn}$. For a complete proof to show that the diffusion part of this renormalized price process satisfies a weak convergence in the Skorokhod topology, see Swishchuk and Vadori, (2017). In applications, such as algorithmic and HFT, and in LOB modeling (when transactions happen every milliseconds), one usually considers the long scale $tn$ (minutes, hours, etc.) instead of $t$ (milliseconds) to have time to prepare solutions for trading problems such as liquidation, acquisitions, and market making, to name a few, where in this paper we will study the first two. In this case, the limit theorems that use the law of large numbers (averaging) for the case where $n$ goes to infinity and $t$ is fixed are very useful.
Following along in a similar fashion, $\varsigma_{SM}=\frac{\sigma^{*}}{\sqrt{\tau}}$ is another new coefficient representing the Semi-Markov diffusion approximation for the jump part of Equation (2). Here, $\sigma^{*}$ is a constant depending on the ergodic and transition probabilities of the Markov chain $X_{k}$ and $\tau$ refers to the interarrival time of the jumps. Then for our two state Markov chain, $\sigma^{*}$ can be defined, as in Swishchuk and Vadori, (2017), as,
$(\sigma^{*})^{2}=\sqrt{4\delta^{2}\left(\frac{1-p_{cont}^{\text{'}}+\pi^{*}(p_{cont}^{\text{'}}-p_{cont})}{(p_{cont}+p_{cont}^{\text{'}}-2)^{2}}\right)},$ (5)
@@ 2.2 Jumps as a function of a Hawkes Process
where $g$ and the diffusion term are as defined before in Section 2.1. In Equation (8), $a$ maps the current state to a price movement, $X_{k}$ is again an ergodic Markov chain with ergodic probabilities $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$, where the current state is computed using information from the previous state. Here, $N(t)$ is again a counting process independent of $X_{k}$ and $W_{t}$. Then using the diffusion approximation for the jump part in Swishchuk and Huffman, (2020), this becomes,
$\begin{split}S_{t}&=S_{0}\pm(g(\nu_{t})\eta_{HP})t+\sqrt{\sigma^{2}+\bar{\sigma}^{2}_{HP}+\left(\sigma^{*}\sqrt{\frac{\lambda}{1-\hat{\mu}}}\right)^{2}}W_{t},\ S_{0}=S\\ &=S_{0}\pm(g(\nu_{t})\eta_{HP})t+\sqrt{\sigma^{2}+\bar{\sigma}^{2}_{HP}+\varsigma_{HP}^{2}}W_{t},\ S_{0}=S.\end{split}$ (9)
The scaling limit for the diffusion approximation part of the midprice process in Equation (9) is formulated via the functional central limit theorem (FCLT) for the General Compound Hawkes process in Swishchuk and Huffman, (2020), where as before in Section 2.1, we use the scaling limit, $\lim_{n\to\infty}$, for $S_{tn}$, which holds in a weak sense in the Skorokhod topology. A proof of this claim can be found in Swishchuk and Huffman, (2020). Similarly to the argument given in Section 2.1, the usefulness in Equation (9) is in terms of being able to prepare solutions to algorithmic and HFT trading problems, where the FCLT enables the use of a diffusion approximation, for the case when $n$ goes to infinity and $t$ is fixed. Here, the drift coefficient can be defined, as in Roldan Contreras, (2023), as,
@@
where $a^{*}$ is a constant depending on the transition probabilities of the Markov chain. This can be defined, as in Swishchuk and Huffman, (2020), as $a^{*}:=\sum_{i\in X}\pi_{i}^{*}a(i)$, where $\pi^{*}$ are the ergodic probabilities of the Markov chain $X$. $\lambda$ is again the background intensity and $\mu$ refers to the output of the response/excitation function. Here, $\varsigma_{HP}=\sigma^{*}\sqrt{\frac{\lambda}{1-\hat{\mu}}}$. $\sigma^{*}$ is again a constant depending on the ergodic probabilities of the Markov chain $X_{k}$. Since this kind of Markov chain has two states, where $X_{k}\in\{-\delta,\delta\}$, $\sigma^{*}$ can be defined, as shown by corollary 3 in Swishchuk and Huffman, (2020), as